OpenStudy (anonymous):

solve the following equation: sin(x)+square root of 2=-sin(x)

OpenStudy (foolaroundmath):

\(\sin{x} + \sqrt{2} = -\sin{x}\) \(\implies 2\sin{x} = -\sqrt{2}\) \(\displaystyle \implies \sin{x} = \frac{1}{\sqrt{2}}\) So, \(\displaystyle x = n\pi+(-1)^{n}\frac{\pi}{4}\), where \(n \in \bf{Z}\)